A patient receiving blood transfusion with a bottle containing blood which is elevated 75.0cm above the arm of the patient. If force of 50N moves the blood to the middle which is 4.0 cm long find a diameter of the needle


1
Expert's answer
2021-09-10T12:18:08-0400

The pressure is equal here:


"p_1=p_2,\\
p_1=\rho g (H+h),\\\space\\
p_2=\frac{F}{A}=\frac{4F}{\pi d^2}.\\\space\\

\rho g (H+h)=\frac{4F}{\pi d^2},\\\space\\

d=\frac{4F}{\pi\rho g(H+h)}=\\\space\\
=\frac{4·50}{\pi·1000· 9.8(0.75+0.04)}=8.22·10^{-3}\text{ m}."